Prev Next

SW&IT Meeting Minutes 2020-11-30

Attendees: Carl, David, Igor, Matthieu, Paul, Stephen, Trey, Xianyu, Yang


Labor Reporting

Report auto-generated every Monday at 6:30am, Do not edit this table, Update your entries in smartsheets.

LR20201116 - Bec, Matthieu 64.0
LR20201116 - Holmberg, Carl 64.0
LR20201116 - Edmeades, David 64.0
LR20201116 - Sola, Igor 20.0
LR20201116 - Grenz, Paul 4.0
LR20201116 - Hooper, Stephen 80.0
LR20201116 - Bryant, Trey 40.0
LR20201116 - Zhang, Yang 64.0
LR20201116 - Zhang, Xianyu 40.0

News from management

Previous Action Items

Owner Action Due Date Completion Date

Open topics

  • Upcoming E/D times - obs5 readiness?
    • All night Nov 30
    • All night Dec 1
    • First half Dec 2
  • robs slowdowns, java VIRT
  • LUCI $HOME/.luci auto-cleanup
  • synology status (email alerts)

Team reports


  • Last week
    • Setup the lab enviroment for flashing the DSP firmware with Juan Pablo and Guido
    • Flash the four left and four right DSP firmware with Juan Pablo and Guido
    • Change ASM Elevation Limits & Configuration from AOS dicussion and document
    • Implementation on AOS side for elevation limit change
    • ATT AdSec
    • AO operation support
  • This week
    • ATT AdSec
    • ATT SW
    • Change ASM Elevation Limits & Configuration demonstration and dicussion
    • Implementation on ao-suervisor side for elevation limit change
    • Completed UA harassment training
    • AO operation support


  • Last week
    • OC, Management
    • PP2021
    • ATT SW
    • DMS AO discussion with SOUL team
    • iLocater fiber discussions
  • This week
    • ATT SW - must complete
    • iLocater fiber discussions
    • Argos cal-unit in TCS mini-review


  • Last week:
    • AGW: Email to Dan Cox for guidance on approach to test firmware Q40, Q41, and Q150 registers when modified by oac commands.
    • AHS: Added collection of _STATE parameter to validate reading Cryomech ADAM data from LBTI INDI service.
    • Hardware: Dell workstation, replaced Dell power supplies, PEWIN32 Lenovo laptop, oac-dev laptop, mock UMAC returned to SO/580F.
  • This week:
    • Admin: closeout activities.
    • statserv: Add LBTI and ARGOS overall status updates.


  • Previous:
    • Mainly TMS work (SPIE paper, software tweak and wiki edits).
  • This week:
    • Plan to focus more on Cravat/OSCO and make some progress.
    • Some misc LBC and MODS stuffs.


  • More phishing stuff; no indications of problems
  • VPN accounts
  • Note new entry into remote observing spreadsheet; at what point is the decision made for these to get VPN access?
  • Office visit this week, planning Thursday evening
    • UPS battery work
    • Drive swap unless Stephen gets to it first
  • UA infosec training


  • domain renewal --
  • issues with non-SSL
  • Wiki reset for Andrew
  • LBC FPIA monitor on obs1 for Steve A.
  • Robs stuff
  • PEPSI HV with Leroy, vm2.mountain rebirth
  • PEPSI reconditioned machines (network)
  • UITS switches, MACs, required training, etc.
  • PCIe card for DIMM, RFQ for CameraLink devices


  • submit final MAC addresses
  • mountain work for second robs and vm2 HV.
  • cleanup of info pages and things
  • grouper stuff with UITS


Last week:
  • LUCI
    • RPMs generation.
    • Continue Geirs script test Barry's night log.

This week:
  • Harassment training
  • LUCI
    • RPM
    • IT 8231 GEIRS timeout
    • IT 8220
    • split current luci motor privilege into two to allow more control on what people can and can't do.
    • Look into reading WebIO status and passing it to the database/telemetry.
    • Add more debugging information in Journalizer DB storing logic. Trying to figure out why sometimes the mask config file doesn't upload.
  • PIT
    • pit admin tool

Topic revision: r14 - 07 Dec 2020, MatthieuBec
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback